OpenStudy (anonymous):

Evaluate the expression, reducing to simplest form and without using a calculator.

OpenStudy (anonymous):

\[\log_{3} \sqrt{243}\]

OpenStudy (anonymous):

that's the question.. i got up to this: \[0.5\log_{3}243 \]

OpenStudy (anonymous):

but i don't know where to go from there.

OpenStudy (anonymous):

\[ \log_{3} \sqrt{243} = \log_{3} \sqrt{3^5} = \frac 5 2 \log_3 3 =\frac 5 2 \]

OpenStudy (anonymous):

ohh. damn. so i was approaching this completely wrong. thank you! ^^

OpenStudy (anonymous):

Glad to help :)

OpenStudy (anonymous):

i had one more question.. could you help? I need to make this all one log function: 3logx−2log(x^2+1)+2log(x−1) i got: (x^3)/(x^2+1)^2*(x-1)^2 but that's wrong. ><

OpenStudy (anonymous):

Sure, but please post it as a new thread.